OpenStudy (anonymous):

How can i find out the LCM of 6, 12, 15?

OpenStudy (anonymous):

you have to think of a number that is able to go into 6, 12, and 15

OpenStudy (anonymous):

for example, does 2 work? 6/2 = 3 so 2 works with 6 12-2=6 so 2 works with 12 15/2=7.5 so 2 is not a factor of 15 therefore you need to find another number

OpenStudy (anonymous):

then try 3, i know 3 is the answer because 3 is able to go evenly into all three numbers

OpenStudy (lgbasallote):

basically it's trial and error but there are methods like prime factorization, decomposition method (i think that's what it was called), etc.

OpenStudy (anonymous):

ope, I don't remember that, I can just look at those numbers and say that 3 is the answer

OpenStudy (anonymous):

I need to find the answer not the factor.

OpenStudy (anonymous):

oh shoottttt, my bad, I read it wrong, I thought it said LCD.....awkkk